OpenStudy (anonymous):

When converting Riemann Sums to Definite Integrals, under what conditions are you allowed to change the limits of integration? I don't mean changing it from 0 to 1 to like 0 to 5, I mean like 0 to 1 to like 1 to 2.

OpenStudy (anonymous):

You can always change the limits of integration. The summation index is just a dummy variable.

OpenStudy (anonymous):

dummy number*

OpenStudy (anonymous):

This assumes you shift all of the terms in the summation by the same amount though.

OpenStudy (anonymous):

So, if f(x) is what you're integrating, and f(x) = x^3, then changing the limits wouldn't be the same?

OpenStudy (anonymous):

Because I've seen cases where it doesn't equate to the same thing, just trying to find some examples right now.

OpenStudy (anonymous):

You can change the limits. Just make sure you shift x^3 byt the same amount.

OpenStudy (anonymous):

Here i'll post this.

OpenStudy (anonymous):

Go a bit down to shifting indices.

OpenStudy (anonymous):

Hmm, I checked Paul's notes, but not quite sure if Power Series and Riemann sums are quite the same. But at the example, would this be correct? \[\lim_{n \rightarrow \infty}\sum_{i = 1}^{n}\frac1n\left(1 + \frac in\right)^{3} = \int_1^2x^3dx\]

OpenStudy (anonymous):

No they're not the same thing but the logic is the same.

OpenStudy (anonymous):

Or would it be \[\lim_{n \rightarrow \infty}\sum_{i = 1}^{n}\frac1n\left(1 + \frac in\right)^{3} = \int_0^1(x + 1)^3dx\]?

OpenStudy (anonymous):

The latter I believe.

OpenStudy (anonymous):

Ok, that's what I thought, but there are cases where I've seen the first one used, and I always end up guessing when picking one or the other. How do I know which form to use, specifically the constant?

OpenStudy (anonymous):

I'm going to flat out say I can't answer this part. Sorry :/ . No idea.

OpenStudy (anonymous):

Alright, no problem :)

OpenStudy (anonymous):

Well, unless someone else can give me an answer, I'm going to assume that changing the limits of integration is only when the constant has literally no relationship to the actual variables. Like \(\lim_{n \rightarrow \infty}\sum_{i = 1}^n\frac 1n((\frac in + 1)^3 + 4) = \int_4^5(x + 1)^3dx\)